OpenStudy (anonymous):

The solution set of sin^2n (theta) + cosec^2n(theta)=1?

OpenStudy (ujjwal):

Is it: 1) \(sin^2n\theta\) or 2) \(\sin^{2n}\theta\) which one 1 or 2?

OpenStudy (anonymous):

2 nd one

OpenStudy (ujjwal):

\[\sin^{2n}\theta+\frac{1}{sin^{2n}\theta}=1\]\[\frac{\sin^{4n}\theta+1}{sin^{2n}\theta}=1\]\[sin^{4n}\theta-sin^{2n}\theta+1=0\] That doesn't give real solution.